OpenStudy (izzy.jane.rasmussen.hairlfip..):

PLEASE HELP!!! The length of a rectangle is 2 cm more than four times the width. If the perimeter of the rectangle is 84 cm, what are its dimensions? Show your work!

OpenStudy (freckles):

let's call the length of the rectangle L and let's call the width of the rectangle W can you write an equation from the first sentence ? It says "L is 2 cm more than 4 times W "

OpenStudy (izzy.jane.rasmussen.hairlfip..):

hmmmmm i think just a second

OpenStudy (izzy.jane.rasmussen.hairlfip..):

would it be 4w+2=L

OpenStudy (freckles):

awesome!!!!

OpenStudy (freckles):

ok let's go to the second sentence/question

OpenStudy (izzy.jane.rasmussen.hairlfip..):

ok!

OpenStudy (freckles):

the perimeter of the rectangle is 84 cm

OpenStudy (freckles):

The perimeter of a rectangle can be found by adding up all the sides

OpenStudy (freckles):

2 sides will have length L and 2 sides will have length W

OpenStudy (izzy.jane.rasmussen.hairlfip..):

and it will all have to add up to 84 right?

OpenStudy (freckles):

so perimeter of a rectangle can be found by calculating 2L+2W right so we also have the equation 2L+2W=84

OpenStudy (freckles):

We have a system of equations to solve: 4W+2=L 2L+2W=84 You can do a substitution of the first equation into the second.

OpenStudy (freckles):

you see the first equation is already solved for L just replace the L in the next equation with what it is equal to

OpenStudy (freckles):

and solve for W first and we will come back and solve for L

OpenStudy (izzy.jane.rasmussen.hairlfip..):

ok so just a minute

OpenStudy (freckles):

k

OpenStudy (izzy.jane.rasmussen.hairlfip..):

yeah im not at all sure how to solve for w.....

OpenStudy (izzy.jane.rasmussen.hairlfip..):

i thought i did but nope i was really wrong

OpenStudy (freckles):

4W+2=L 2L+2W=84 See that the first equation is L=4W+2 just replace the L in 2L+2W=84 with 4W+2 since you do have L=4W+2 so you have 2(4W+2)+2W=84 this is a linear equation in terms of one variable

OpenStudy (freckles):

First step distribute where you have that 2(4W+2) distributive law is a(b+c)=a*b+a*c

OpenStudy (izzy.jane.rasmussen.hairlfip..):

ok? give me a minute

OpenStudy (freckles):

a(b+c)=a*b+a*c 2(4W+2)=2*4W+2*2=....

OpenStudy (izzy.jane.rasmussen.hairlfip..):

ok that makes more sense give me a minute

OpenStudy (izzy.jane.rasmussen.hairlfip..):

nope nevermind got something weird idk what i did wrong but i dont think its w = 7

OpenStudy (freckles):

2*4W+2*2=8W+4 did you get this part?

OpenStudy (freckles):

\[2L+2W=84 \text{ where } L=4W+2 \\ 2(4W+2)+2W=84 \\ \\ 8W+4+2W=84 \\ \text{ now combine like terms on the left hand side }\]

OpenStudy (izzy.jane.rasmussen.hairlfip..):

uhh yes i did ok um i actually have to go thank you im going to copy this and use it for reference for tonight when i finish this thank you for all your help

OpenStudy (freckles):

Your W was just a little off I bet you will get it have a good day

OpenStudy (izzy.jane.rasmussen.hairlfip..):

you too!